OpenStudy (anonymous):

Simplify (5x+10)(5x-10)

OpenStudy (anonymous):

FOIL

OpenStudy (anonymous):

so...answer would be 5(x^2-4)?

OpenStudy (dumbcow):

factor out a 5 5(x+2)(x-2)

OpenStudy (anonymous):

This is simple, you can solve by using the identity: (a + b) (a-b) = a^2 - b^2 so it is (5x)^2 - (10)^2 = 25x^2 - 100

OpenStudy (dumbcow):

ooops yeah i meant 25(x-2)(x+2)

OpenStudy (anonymous):

which reduces to 5(x^2-4)?

OpenStudy (anonymous):

Then you can take out common factor 25 to get final answer 25 (x^2 - 4)

OpenStudy (anonymous):

@kaynkay Go thru my answer. It is step-by-step logic and clear

OpenStudy (anonymous):

oops - I see, have to leave the 25, since the 100 is the factor - so answer would be: 25(x^2-4)...got it, thank you!

OpenStudy (anonymous):

Good ! However, if you hv to show factors, then dumbcow's answer is right.